OpenStudy (anonymous):

What is the y intercept of -6x -1/3y = 6? Will Medal!

OpenStudy (anonymous):

subtitute 0 for x and solve: -6*0 - 1/3y=6 -1/3y=6 multiply each side by -3 y=-18

OpenStudy (anonymous):

oh ok thanks so much! that makes perfect sense!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

could you answer another problem I have? @ksanka